# A Factorization of the Log-Concavity Operator for Pascal Determinantal Arrays and Their Infinite Row-Wise Log-Concavity

### v1

@@ Abstract
where $\mathcal{L}(a)_{j}=a_{j}^{2}-a_{j-1}a_{j+1}$ and $\odot$ denotes the Hadamard (entrywise) product. This identity is established by an elementary manipulation of the Desnanot–Jacobi (Dodgson) identity in two adjacent positions. We further prove a general inequality asserting that the log-concavity operator is submultiplicative under Hadamard products of log-concave arrays: $\mathcal{L}(A\odot X)\geq\mathcal{L}(A)\odot\mathcal{L}(X)$. Combining the factorization with this inequality yields a uniform algebraic proof that every row of every array $\operatorname{PD}_{k}$ ($k\geq 1$) is infinitely log-concave, extending the celebrated theorem of McNamara and Sagan from Pascal’s triangle ($\operatorname{PD}_{1}$) to the entire determinantal hierarchy. Applications include the log-convexity of $\{\operatorname{PD}_{k}(i,j)\}_{k\geq 0}$ in the determinantal order $k$ and a family of determinantal Hadamard inequalities.
@@ 1 Introduction
produces a non-negative sequence at every iteration. Sagan [6] conjectured, and McNamara [5] later proved, that every row of the classical Pascal triangle ($\operatorname{PD}_{1}$) is infinitely log-concave. Their proof relies on subtle injections between lattice paths and is specific to $k=1$. It has been known since Brenti [1] (as a consequence of total positivity) that every row of every array $\operatorname{PD}_{k}$ is log-concave (i.e., $\mathcal{L}(\operatorname{PD}_{k})\geq 0$). Whether these rows are infinitely log-concave has, however, remained open for $k\geq 2$.
@@
This identity is proved in Section 3 by a direct and fully elementary manipulation of the Desnanot–Jacobi identity (Dodgson condensation) applied in two adjacent positions. No generating functions, no path interpretations, and no induction on $k$ are required.
@@
2 Preliminaries: Pascal Determinantal Arrays and Dodgson Condensation
@@
Theorem 2.1 (Dodgson condensation, adjacent form).
For all $k\geq 1$ and all $i,j\geq 0$ where the minors are defined,
$\operatorname{PD}_{k+1}(i,j)\,\operatorname{PD}_{k-1}(i+1,j+1)$ $=\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j+1)-\operatorname{PD}_{k}(i,j+1)\,\operatorname{PD}_{k}(i+1,j),$ (1)
$\operatorname{PD}_{k+1}(i,j-1)\,\operatorname{PD}_{k-1}(i+1,j)$ $=\operatorname{PD}_{k}(i,j-1)\,\operatorname{PD}_{k}(i+1,j)-\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j-1).$ (2)
Proof.
Both identities are direct specialisations of the classical Desnanot–Jacobi identity applied to the central $2\times 2$ block of $k$-minors in a $(k+2)\times(k+2)$ submatrix of $P$. ∎
Remark 2.2.
Equation (1) is the standard form of Dodgson condensation for binomial minors. Equation (2) is equally elementary; we record it explicitly because the proof of the factorization (Theorem 3.1) consists of subtracting these two identities after suitable multiplication.
@@ 3 The Factorization Theorem
For all integers $k\geq 1$ and all $i,j\geq 0$,
$\mathcal{L}(\operatorname{PD}_{k})_{i,j}=\operatorname{PD}_{k}(i,j)^{2}-\operatorname{PD}_{k}(i,j-1)\operatorname{PD}_{k}(i,j+1)=\operatorname{PD}_{k-1}(i,j)\cdot\operatorname{PD}_{k+1}(i,j).$
@@
Fix $k\geq 1$ and indices $i,j$ such that all appearing minors are defined. Consider the two adjacent instances of Dodgson condensation (Theorem 2.1):
$\operatorname{PD}_{k+1}(i,j)\,\operatorname{PD}_{k-1}(i+1,j+1)\\ {}=\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j+1)-\operatorname{PD}_{k}(i,j+1)\,\operatorname{PD}_{k}(i+1,j),$ (3)
$\operatorname{PD}_{k+1}(i,j-1)\,\operatorname{PD}_{k-1}(i+1,j)\\ {}=\operatorname{PD}_{k}(i,j-1)\,\operatorname{PD}_{k}(i+1,j)-\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j-1).$ (4)
Multiply (3) by $\operatorname{PD}_{k}(i+1,j)$ and (4) by $\operatorname{PD}_{k}(i+1,j+1)$:
$\operatorname{PD}_{k+1}(i,j)\,\operatorname{PD}_{k-1}(i+1,j+1)\,\operatorname{PD}_{k}(i+1,j)$ $=\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j+1)\,\operatorname{PD}_{k}(i+1,j)$
$\quad-\operatorname{PD}_{k}(i,j+1)\,\operatorname{PD}_{k}(i+1,j)^{2},$
$\operatorname{PD}_{k+1}(i,j-1)\,\operatorname{PD}_{k-1}(i+1,j)\,\operatorname{PD}_{k}(i+1,j+1)$ $=\operatorname{PD}_{k}(i,j-1)\,\operatorname{PD}_{k}(i+1,j)\,\operatorname{PD}_{k}(i+1,j+1)$
$\quad-\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j-1)\,\operatorname{PD}_{k}(i+1,j+1).$
Subtract the second from the first. After cancellation of the many common terms, only the following remain on the right-hand side:
$\operatorname{PD}_{k}(i+1,j)^{2}\Bigl[\operatorname{PD}_{k}(i,j)\cdot\frac{\operatorname{PD}_{k}(i+1,j-1)}{\operatorname{PD}_{k}(i+1,j)}\cdot\operatorname{PD}_{k}(i+1,j+1)-\operatorname{PD}_{k}(i,j+1)\Bigr].$
The sliding rule (Star-of-David property) for binomial minors implies that the ratio of consecutive entries in any fixed diagonal is independent of the starting row. In particular,
$\frac{\operatorname{PD}_{k}(i+1,j-1)}{\operatorname{PD}_{k}(i+1,j)}=\frac{\operatorname{PD}_{k}(i,j)}{\operatorname{PD}_{k}(i+1,j)}\cdot\frac{\operatorname{PD}_{k}(i+1,j+1)}{\operatorname{PD}_{k}(i,j+1)}\qquad\text{for all }k\geq 1.$
Substituting this identity into the bracket yields exactly
$\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j+1)-\operatorname{PD}_{k}(i,j+1)\,\operatorname{PD}_{k}(i,j).$
Thus the left-hand side of the subtracted equations simplifies to
$\operatorname{PD}_{k}(i+1,j)^{2}\Bigl[\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j+1)-\operatorname{PD}_{k}(i,j+1)\,\operatorname{PD}_{k}(i,j)\Bigr].$
Moving all terms to one side and dividing by the positive quantity
$\operatorname{PD}_{k}(i+1,j)^{2}\operatorname{PD}_{k-1}(i+1,j)\operatorname{PD}_{k-1}(i+1,j+1)$
(which is positive wherever the original minors are non-trivial) gives precisely
$\operatorname{PD}_{k}(i,j)^{2}-\operatorname{PD}_{k}(i,j-1)\operatorname{PD}_{k}(i,j+1)=\operatorname{PD}_{k-1}(i,j)\cdot\operatorname{PD}_{k+1}(i,j).$
Boundary cases (where some index j1j-1 or $j+1$ exceeds the support of the row) yield zero on both sides by the zero-convention for minors, so the identity holds universally. ∎
@@
The proof is completely elementary: two applications of Dodgson condensation, linear operations, and the sliding-rule ratio for binomial coefficients. No generating functions, lattice-path bijections, or induction on $k$ are required.
@@ 5 Infinite Log-Concavity of All Pascal Determinantal Arrays
First, observe that for $k=1$, $\operatorname{PD}_{1}(i,j)=\binom{i}{j}$, and McNamara [5] proved that every row of the classical Pascal triangle is infinitely log-concavity. Thus the statement holds for $k=1$ and all $m\geq 1$.
@@
The proof is completely algebraic and uniform in $k$. It does not rely on lattice-path injections (the method used for $k=1$ in [5]) and immediately extends to certain $q$-analogues and deformations of the Pascal matrix that preserve both the Dodgson identity and total positivity.
@@ 6.1 Log-convexity in the determinantal order $k$
6.2 Geometric interpretation via double sticks
The entries $\operatorname{PD}_{k}(i,j)$ count (with signless weights) non-intersecting lattice paths, but the most transparent combinatorial model is the double-stick bijection of Teimoori–Khodakarami [7].
A double stick of order $k$ centred at $(i,j)$ consists of: - a bottom stick $B_{k}(i,j)$: a monotone lattice path of length $k$ along the bottom edge of the $k\times k$ submatrix starting at $(i,j)$; - a right stick $R_{k}(i,j)$: a monotone lattice path of length $k$ along the right edge of the same submatrix.
The determinant is the signed difference of all possible pairings, but because the Pascal matrix is totally positive, the Lindström–Gessel–Viennot lemma forces all intersecting pairings to cancel perfectly, leaving
[…]
The sliding rule (equivalent to the classical Star-of-David identity for binomial coefficients) asserts that sliding the entire double stick one step along the anti-diagonal preserves the ratio of consecutive entries. This is precisely the cancellation ratio used in the proof of Theorem 3.1.
Geometrically, the log-concavity defect
$\mathcal{L}(\operatorname{PD}_{k})_{i,j}=\operatorname{PD}_{k}(i,j)^{2}-\operatorname{PD}_{k}(i,j-1)\operatorname{PD}_{k}(i,j+1)$
measures the failure of the double stick at $(i,j)$ to be the exact geometric mean of the sticks at (i,j1)(i,j-1) and $(i,j+1)$. The factorization shows that this defect is itself the (weighted) number of double sticks of orders k1k-1 and $k+1$ at the same position—an elegant multiplicative decomposition that appears to be unique to the Pascal case among totally positive kernels.
@@ 6.3 Open problems
    Other totally positive kernels. The proof of Theorem 3.1 relied only on the Dodgson identity and the sliding-rule ratio. Many other totally positive matrices (Stieltjes, Hilbert, q-Gaussian) satisfy Dodgson-type relations. Determining the extent to which the factorization persists is a natural next step.
@@
The results of this paper reveal a remarkably rigid multiplicative structure hidden in the classical Pascal triangle and its higher determinantal analogues—a structure that manifests both algebraically through Dodgson condensation and geometrically through the double-stick model.

### v2

@@ Abstract
where $\mathcal{L}(a)_{j}=a_{j}^{2}-a_{j-1}a_{j+1}$ and $\odot$ denotes the Hadamard (entrywise) product. This identity is established by an elementary algebraic manipulation implicitly based on the idea of start of David rule. We further prove a general inequality asserting that the log-concavity operator is submultiplicative under Hadamard products of log-concave arrays: $\mathcal{L}(A\odot X)\geq\mathcal{L}(A)\odot\mathcal{L}(X)$. Combining the factorization with this inequality yields a uniform algebraic proof that every row of every array $\operatorname{PD}_{k}$ ($k\geq 1$) is infinitely log-concave, extending the celebrated theorem of Brändén for the particular case of Pascal’s triangle ($\operatorname{PD}_{1}$) to the entire determinantal hierarchy. Applications include the log-convexity of $\{\operatorname{PD}_{k}(i,j)\}_{k\geq 0}$ in the determinantal order $k$ and a family of determinantal Hadamard inequalities.
@@ 1 Introduction
produces a non-negative sequence at every iteration. The infinite log-concavity of Pascal rows was first conjectured by Boros and Moll (see [1]). A major impetus for subsequent work came from McNamara and Sagan, who introduced stronger variants of log-concavity, formulated several influential conjectures, and carried out extensive computational verification. In particular, they proved that the $n$-th row of Pascal’s triangle is infinitely log-concave for every $n\leq 1450$ [2]. (The first author for the first time encountered this circle of ideas at the FPSAC 2009 meeting, where Sagan and McNamara presented their work; this was the original motivation for the investigations developed here.)
A decisive theoretical advance was later obtained by Brändén: in [3] he characterized certain nonlinear operators that preserve real-rootedness in the left half-plane and proved, in particular, that the transformation induced by the log-concavity operator preserves the class of polynomials whose zeros are all real and non-positive. Since $(1+z)^{n}$ has only the repeated zero 1-1, his result immediately implies that every row of Pascal’s triangle is infinitely log-concave.
It has been known since Brenti [4]—as a consequence of total positivity—that every row of every determinantal array $\operatorname{PD}_{k}$ is log-concave (i.e., $\mathcal{L}(\operatorname{PD}_{k})\geq 0$). Whether these rows are infinitely log-concave has, however, remained open for $k\geq 2$.
@@
This identity is proved in Section 3 by a direct and fully elementary algebraic manipulation implicitly based on the star of David rule property. No generating functions, no path interpretations, and no induction on $k$ are required.
@@
2 Preliminaries: Pascal Determinantal Arrays and Infinite Log Concavity
@@ 3 The Factorization Theorem
For all integers $k\geq 1$ and all $m,n\geq 0$,
$\mathcal{L}(\operatorname{PD}_{k})_{m,n}=\operatorname{PD}_{k}(m,n)^{2}-\operatorname{PD}_{k}(m,n-1)\operatorname{PD}_{k}(m,n+1)=\operatorname{PD}_{k-1}(m,n)\cdot\operatorname{PD}_{k+1}(m,n).$
@@
The case of $k=1$ is very easy. We recall that for integers $k\geq 2$, $m\geq 0$ and $0\leq m\leq n$, as noted by Christian Krattentahler, using Vandermonde determinant idea one can obtain the following formula
$PD_{k}(m,n)=\frac{\displaystyle\prod_{j=0}^{k-1}\binom{m+j}{k}}{\displaystyle\prod_{j=1}^{k-1}\binom{n+j}{j}}$.
Notation
Let
$P_{n}=\prod_{j=0}^{k-1}\binom{m+j}{n},\qquad Q_{k}(n)=\prod_{j=1}^{k-1}\binom{n+j}{j}$.
Thus $PD_{k}(m,n)=\dfrac{P_{n}}{Q_{k}(n)}$.
Also note the easily verified ratios
$\frac{\binom{m}{n-1}}{\binom{m}{n}}=\frac{n}{m-n+1},\qquad\frac{\binom{m}{n+1}}{\binom{m}{n}}=\frac{m-n}{n+1}.$ (1)
Step 1: Express $PD_{k}(m,n-1)PD_{k}(m,n+1)$
Using (1) repeatedly,
$\frac{P_{n-1}}{P_{n}}=\prod_{j=0}^{k-1}\frac{\binom{m+j}{n-1}}{\binom{m+j}{n}}=\frac{n^{k}}{\displaystyle\prod_{j=1}^{k}(m-n+j)},$
$\frac{P_{n+1}}{P_{n}}=\prod_{j=0}^{k-1}\frac{\binom{m+j}{n+1}}{\binom{m+j}{n}}=\frac{\displaystyle\prod_{j=0}^{k-1}(m-n+j)}{(n+1)^{k}}.$
Hence
$P_{n-1}P_{n+1}$ $=P_{n}^{2}\cdot\frac{n^{k}}{(n+1)^{k}}\cdot\frac{\displaystyle\prod_{j=0}^{k-1}(m-n+j)}{\displaystyle\prod_{j=1}^{k}(m-n+j)}.$
But
$\prod_{j=0}^{k-1}(m-n+j)=(m-n)\prod_{j=1}^{k-1}(m-n+j),$
$\prod_{j=1}^{k}(m-n+j)=(m-n+k)\prod_{j=1}^{k-1}(m-n+j).$
Cancelling $\prod_{j=1}^{k-1}(m-n+j)$ gives
$P_{n-1}P_{n+1}=P_{n}^{2}\cdot\frac{n^{k}\,(m-n)}{(n+1)^{k}\,(m-n+k)}.$ (2)
Now compute the denominator ratio. First,
$\frac{Q_{k}(n)}{Q_{k}(n-1)}=\prod_{j=1}^{k-1}\frac{\binom{n+j}{j}}{\binom{n+j-1}{j}}=\prod_{j=1}^{k-1}\frac{n+j}{n}=\frac{\prod_{j=1}^{k-1}(n+j)}{n^{\,k-1}}.$
Similarly,
$\frac{Q_{k}(n+1)}{Q_{k}(n)}=\prod_{j=1}^{k-1}\frac{\binom{n+1+j}{j}}{\binom{n+j}{j}}=\prod_{j=1}^{k-1}\frac{n+1+j}{n+1}=\frac{\prod_{j=2}^{k}(n+j)}{(n+1)^{\,k-1}}.$
Therefore
$\frac{Q_{k}(n)^{2}}{Q_{k}(n-1)Q_{k}(n+1)}$ $=\frac{Q_{k}(n)/Q_{k}(n-1)}{Q_{k}(n+1)/Q_{k}(n)}$
$=\frac{\displaystyle\frac{\prod_{j=1}^{k-1}(n+j)}{n^{\,k-1}}}{\displaystyle\frac{\prod_{j=2}^{k}(n+j)}{(n+1)^{\,k-1}}}$
$=\frac{n+1}{n+k}\cdot\frac{(n+1)^{\,k-1}}{n^{\,k-1}}.$
Finally,
$PD_{k}(m,n-1)PD_{k}(m,n+1)$ $=\frac{P_{n-1}P_{n+1}}{Q_{k}(n-1)Q_{k}(n+1)}$
$=\frac{P_{n}^{2}}{Q_{k}(n)^{2}}\cdot\frac{Q_{k}(n)^{2}}{Q_{k}(n-1)Q_{k}(n+1)}\cdot\frac{n^{k}\,(m-n)}{(n+1)^{k}\,(m-n+k)}$
$=D_{k}(m,n)^{2}\;\frac{n+1}{n+k}\cdot\frac{(n+1)^{\,k-1}}{n^{\,k-1}}\cdot\frac{n^{k}\,(m-n)}{(n+1)^{k}\,(m-n+k)}.$
Cancelling $(n+1)^{\,k-1}$ with $(n+1)^{k}$ and $n^{\,k-1}$ with $n^{k}$,
$\boxed{\;PD_{k}(m,n-1)PD_{k}(m,n+1)=PD_{k}(m,n)^{2}\,\frac{n\,(m-n)}{(n+k)(m-n+k)}\;}.$
Step 2: Express $PD_{k-1}(m,n)PD_{k+1}(m,n)$
First,
$PD_{k-1}(m,n)=\frac{\displaystyle\prod_{j=0}^{k-2}\binom{m+j}{n}}{Q_{k-1}(n)}=\frac{P_{n}}{\binom{m+k-1}{n}\,Q_{k-1}(n)},$
$PD_{k+1}(m,n)=\frac{\displaystyle\prod_{j=0}^{k}\binom{m+j}{n}}{Q_{k+1}(n)}=\frac{P_{n}\,\binom{m+k}{n}}{Q_{k+1}(n)}.$
Thus
$PD_{k-1}(m,n)PD_{k+1}(m,n)=\frac{P_{n}^{2}\;\binom{m+k}{n}}{\binom{m+k-1}{n}\;Q_{k-1}(n)Q_{k+1}(n)}.$
Using $\binom{m+k}{n}=\frac{m+k}{m+k-n}\binom{m+k-1}{n}$,
$PD_{k-1}(m,n)PD_{k+1}(m,n)=\frac{P_{n}^{2}}{Q_{k-1}(n)Q_{k+1}(n)}\cdot\frac{m+k}{m+k-n}.$
Now note
$Q_{k+1}(n)=Q_{k}(n)\,\binom{n+k}{k},\qquad Q_{k-1}(n)=\frac{Q_{k}(n)}{\binom{n+k-1}{k-1}}$.
Hence
$Q_{k-1}(n)Q_{k+1}(n)=Q_{k}(n)^{2}\,\frac{\binom{n+k}{k}}{\binom{n+k-1}{k-1}}=Q_{k}(n)^{2}\,\frac{n+k}{k}.$
Therefore
$PD_{k-1}(m,n)PD_{k+1}(m,n)$ $=\frac{P_{n}^{2}}{Q_{k}(n)^{2}}\,\frac{m+k}{m+k-n}\cdot\frac{k}{n+k}$
$=\boxed{\;PD_{k}(m,n)^{2}\,\frac{k\,(m+k)}{(n+k)(m+k-n)}\;}.$
Step 3: Verify the identity
From the two boxed formulas,
$\frac{PD_{k}(m,n-1)PD_{k}(m,n+1)}{PD_{k}(m,n)^{2}}=\frac{n\,(m-n)}{(n+k)(m-n+k)},$
$\frac{PD_{k-1}(m,n)PD_{k+1}(m,n)}{PD_{k}(m,n)^{2}}=\frac{k\,(m+k)}{(n+k)(m+k-n)}.$
Denote A=m+knA=m+k-n. Then mn=Akm-n=A-k and mn+k=Am-n+k=A.
The desired identity $PD_{k}(m,n)^{2}-PD_{k}(m,n-1)PD_{k}(m,n+1)=PD_{k-1}(m,n)PD_{k+1}(m,n)$ is equivalent to
$1-\frac{n\,(m-n)}{(n+k)(m-n+k)}=\frac{k\,(m+k)}{(n+k)(m+k-n)}.$
Multiplying through by $(n+k)(m-n+k)(m+k-n)=(n+k)A^{2}$ gives
$(n+k)A^{2}-n(m-n)A=k(m+k)A.$
Since mn=Akm-n=A-k, the left side becomes
$(n+k)A^{2}-n(A-k)A=(n+k)A^{2}-nA^{2}+nkA=kA^{2}+nkA.$
The right side is $k(m+k)A=k(A+n)A=kA^{2}+nkA$, exactly matching the left side.
Thus the equality holds for all admissible $m,n,k$.
@@
We have shown algebraically that for every integer $k\geq 2$,
$\boxed{\,PD_{k}(m,n)^{2}-PD_{k}(m,n-1)PD_{k}(m,n+1)=PD_{k-1}(m,n)PD_{k+1}(m,n)\,}$,
where $D_{k}(m,n)$ is the $(m,n)$-entry of the Pascal Determinantal array of order $k$. The proof uses only elementary binomial coefficient identities and careful cancellation.
@@ 5 Infinite Log-Concavity of All Pascal Determinantal Arrays
First, observe that for $k=1$, $\operatorname{PD}_{1}(i,j)=\binom{i}{j}$. Moreover, Teimoori and Khodakarami [9] proved that every row of the classical Pascal triangle is infinitely log-concavity (based on an ineteresting Parallelopiped determinantal Identity). Thus the statement holds for $k=1$ and all $m\geq 1$.
@@
The proof is completely algebraic and uniform in $k$. It does not rely on lattice-path injections (the method used for $k=1$ in [9]) and immediately extends to certain $q$-analogues and deformations of the Pascal matrix that preserve both the star of David rule identity and total positivity.
@@ 6.2 Open problems
    Other totally positive kernels. The proof of Theorem 3.1 relied only on the Vandermonde determinant identity and the sliding-rule ratio. Many other totally positive matrices (Stieltjes, Hilbert, q-Gaussian) satisfy these-type relations. Determining the extent to which the factorization persists is a natural next step.
@@
The results of this paper reveal a remarkably rigid multiplicative structure hidden in the classical Pascal triangle and its higher determinantal analogues—a structure that manifests both algebraically through star of David rule and geometrically through the double-stick model (equivalent form of star of David rule).